OpenStudy (surryyy):

Edward worked these hours during the week: 6, 5 1/2, 4 1/4, 7 1/2, 4 3/4. If he is paid $5.15 per h, how much did he earn? Answer $133.90 $141.63 $144.20 $139.05

OpenStudy (anonymous):

Total Hours = 6+5.5+4.25+7.5+4.75 = 28 hours 28*5.15= $144.20

OpenStudy (surryyy):

thnx!

OpenStudy (anonymous):

no prob!